# Hi-Core: Hierarchical Knowledge Transfer for Continual Reinforcement Learning

### v1

@@ Abstract
Continual reinforcement learning (CRL) empowers RL agents with the ability to learn from a sequence of tasks, preserving previous knowledge and leveraging it to facilitate future learning. However, existing methods often focus on transferring low-level knowledge across similar tasks, which neglects the hierarchical structure of human cognitive control, resulting in insufficient knowledge transfer across diverse tasks. To enhance high-level knowledge transfer, we propose a novel framework named Hi-Core (Hierarchical knowledge transfer for Continual reinforcement learning), which is structured in two layers: 1) the high-level policy formulation which utilizes the powerful reasoning ability of the Large Language Model (LLM) to set goals and 2) the low-level policy learning through RL which is oriented by high-level goals. Moreover, the knowledge base (policy library) is constructed to store policies that can be retrieved for hierarchical knowledge transfer. Experiments conducted in MiniGrid have demonstrated the effectiveness of Hi-Core in handling diverse CRL tasks, outperforming popular baselines.
@@ 1 Introduction
Reinforcement learning (RL) is a powerful paradigm in artificial intelligence (AI) that enables agents to learn optimal behaviors through interactions with environments. The capacity to continuously adapt and learn from new tasks without starting from scratch is a defining characteristic of human learning Kudithipudi et al. (2022) and a desirable trait for RL agents deployed in the dynamic and unpredictable real world. To achieve this, the research of continual reinforcement learning (CRL, a.k.a. lifelong reinforcement learning) has emerged Rolnick et al. (2019); Wolczyk et al. (2022); Kessler et al. (2022). It extends traditional RL by empowering agents with the ability to learn from a sequence of tasks, preserving knowledge from previous tasks, and using this knowledge to enhance learning efficiency and performance on future tasks. CRL aims to emulate the human capacity for lifelong learning and represents an ambitious field of research that tackles the challenges of long-term, real-world applications by addressing problems of increasing scale, diversity, and non-stationarity Khetarpal et al. (2022).
@@
Despite progress in CRL, the field continues to face a critical challenge: insufficient transfer of knowledge across diverse tasks. Sufficient knowledge transfer is essential for a robust CRL system, yet many existing methods struggle to achieve this, leading to suboptimal performance and a failure to fully exploit the benefits of continual learning (CL) Wang et al. (2023c). A primary reason for this shortfall is the neglect of the sophisticated hierarchical structure of human cognitive control, which is primarily located in the prefrontal cortex and is complemented by the cerebellum for precise action control, enabling high-level cognition for envisioning the future and planning D’Mello et al. (2020); Friedman and Robbins (2022). This structure enhances humans’ ability to tackle complex tasks, and crucially, transfer skills and high-level knowledge across significantly diverse tasks. In contrast, as shown on the left side of Figure 1, existing CRL methods often focus only on transferring low-level knowledge, such as sharing the policy network. This limits them to sequences of highly similar tasks with slight variations only in goals or parameters Kessler et al. (2022); Gaya et al. (2023), and results in ineffective transfer across diverse tasks.
@@
To take advantage of abstract similarities between diverse tasks, it is essential to represent high-level knowledge effectively and then employ robust models capable of extracting and transferring this knowledge. Human language, with its inherent abstraction, is a natural fit for representing high-level knowledge. Furthermore, recent breakthroughs in large-scale language research have demonstrated the impressive reasoning and in-context learning abilities of the large language model (LLM) Zhao et al. (2023); Xi et al. (2023); Wang et al. (2023b). Based on these insights, we propose a new framework named Hi-Core (Hierarchical knowledge transfer for Continual reinforcement learning), which integrates the powerful capabilities of an LLM into the CRL paradigm to enhance high-level knowledge transfer and eventually improve learning performance.
@@
The right side of Figure 1 illustrates the idea of Hi-Core. Hi-Core is structured in two layers: the high-level policy formulation and the low-level policy learning. At the high level, Hi-Core utilizes the powerful reasoning ability of a LLM to formulate a high-level policy, which is represented as a temporally extended sequence of goals. Each goal consists of a descriptive component and an intrinsic reward function that outlines the intermediate states the agent needs to reach. The high-level policy is evaluated based on feedback from the low-level learning process, allowing the LLM to refine its policy. Following this process, the high-level policy along with the corresponding information is stored in a policy library. Upon encountering a new task, Hi-Core can retrieve this policy from the library as a context of the LLM for enhancing high-level knowledge transfer. At the low level, Hi-Core employs goal-oriented RL to learn a policy that is guided by the formulated high-level policy. The verified policy can be stored in the policy library to reduce the risk of catastrophic forgetting when encountering old tasks. Experiments are conducted in simple yet challenging MiniGrid environments. The results demonstrate that Hi-Core significantly outperforms popular baselines in terms of overall performance and transfer performance. The main contributions of the work are as follows:
We proposed Hi-Core, a novel framework that leverages hierarchical knowledge transfer for CRL, which aligns with human cognitive control, enhancing the effectiveness and generalization of agents.
@@
To the best of our knowledge, Hi-Core is the first work to integrate the powerful reasoning ability of LLM into the CRL paradigm, facilitating high-level policy formulation and knowledge transfer across diverse tasks.
@@
Extensive experiments in MiniGrid provide empirical evidence of Hi-Core’s effectiveness. Specifically, the results indicate that Hi-Core can be applied to heterogeneous tasks, achieving comparable transfer performance.
@@ 2.1 Preliminaries
The continual reinforcement learning process can be formulated as a series of related Markov decision processes (MDPs) . Each MDP represents a different task or problem instance that an agent needs to solve over its lifetime. Here, and denote the state and action space of task , respectively, while is the transition probability function, and is the reward function. At each time step, the learning agent perceives the current state and selects an action according to its policy with parameters . The agent then transitions to the next state and receives a reward . The target of an agent on the task is to maximize the expected return , where is the discount factor of the reward, and is the horizon.
@@ 2.2.1 Continual Reinforcement Learning
Another issue in CRL is transfer learning, which is crucial for efficient policy adaptation. Naive approaches, like fine-tuning, that train a single model on each new task provide good scalability and transferability but suffer from catastrophic forgetting. Regularization-based methods, such as EWC Kirkpatrick et al. (2017); Wang et al. (2023c), have been proposed to prevent this side effect, but often reduce plasticity. Some architectural innovations have been proposed to balance the trade-off between plasticity and stability Rusu et al. (2016); Berseth et al. (2022). Furthermore, methods like OWL Kessler et al. (2022) and MAXQINIT Abel et al. (2018) leverage policy factorization and value function transfer, respectively, for improved learning across tasks.
Most existing methods perform well when applied to sequences of tasks that exhibit high environmental similarity, such as tasks where only specific parameters within the environment are altered or the objectives within the same environment are different. However, their effectiveness is greatly diminished when dealing with a sequence of diverse tasks. Our proposed framework aims to overcome this limitation by leveraging hierarchical knowledge transfer.
@@ 2.2.2 Reinforcement Learning with the LLM
Recent advancements have combined the large language model with reinforcement learning to address reinforcement learning challenges such as sample efficiency and generalization. These integrations utilize LLM’s vast knowledge and reasoning ability to improve agents’ performance. For instance, studies have developed methods for efficient agent-LLM interactions Hu et al. (2023) and unified agent foundations with language models for better experience reuse Palo et al. (2023).
Frameworks that merge LLMs with RL have also emerged, such as LAMP, which pretrains RL agents with Vision-Language Models Adeniji et al. (2023), and HiP, which uses foundation models for long-horizon tasks Ajay et al. (2023). In addition, the RAFA framework offers a principled approach with guarantees to optimize reasoning and actions Liu et al. (2023). Research into LLM-guided skill learning and reward shaping is also gaining traction. The BOSS method Zhang et al. (2023) and the work on LLM-based reward structures for robotics Yu et al. (2023) exemplify this trend, with studies like Kwon et al. (2023) focusing on LLM-driven reward design.
Our proposed framework extends these studies to continual learning in RL by leveraging LLM’s powerful reason ability for high-level policy formulation, thus enhancing high-level knowledge transfer.
@@ 3 Method
In this section, we clarify the research problem and describe the framework.
@@
Problem Formalization. We formalize the CRL problem in the goal-oriented RL framework Chen and Luo (2022); Colas et al. (2022). The agent perceives not only the current state of the environment but is also provided with a specific goal , where represents the space of possible goals. For task and its corresponding goal sequence , the agent’s objective is to learn a policy that effectively maps each state to an appropriate action . The aim is to maximize the expected cumulative reward over time, guided by the sum of the extrinsic reward and the intrinsic rewards Dilokthanakul et al. (2019), where represents the next state in the task . The former is determined by the environment and is usually sparse. The latter is typically defined so that the agent receives a reward if it achieves the given goal.
@@
Overview. Hi-Core is illustrated in Figure 2. For simplicity, we have omitted the task-indexed superscript in the description. The framework orchestrates the interaction between the CRL agent and the environment simulator for each task in the task sequence, aligning with standard RL protocols. For each task, the corresponding environment simulator provides the current state to the agent at each timestep, which then selects an action to perform, resulting in a reward and the next state. Internally, the CRL agent is structured in two layers: the high-level policy formulation and the low-level policy learning. Additionally, a policy library is maintained for the storage and retrieval of policies. The formulation of the high-level policy is initiated when the agent interacts first with the environment. If the policy library is not empty, related experiences (including high-level policies, corresponding environment descriptions and feedback) will be retrieved through the description of the environment to assist in the formulation. Once the high-level policy is established, it guides the learning of the low-level policy. Provided that there is sufficient storage, the most similar low-level policy can be retrieved from the policy library to serve as an advantageous initial point. Through feedback from the RL learning process, the LLM gradually refines its high-level policy to improve performance. Upon completion of a task, the successful policies are stored in the policy library.
3.1 High-level Policy Formulation
@@
To integrate LLM into the goal-oriented RL framework, we conceptualize LLM as a function , where is the set of all prompts, is the set of textual descriptions of the environment, and is the set of textual descriptions of the sequence of goals. The description of the environment can be automatically obtained through a hard-coded reporter Dasgupta et al. (2022). The prompt is a concatenation of two parts: and , where is the system prompt and is the experience retrieved from the policy library. Description of each high-level policy, that is, textual output is mapped to the corresponding goal sequence through a defined parsing function. In order to improve the stability of LLM’s output, we structure the goal into a tuple: , where is the textual description of the goal state, is a verification function , is a intrinsic reward value after achieving the goal. Optional validation functions are provided in .
@@
Intrinsic reward can not only guide the agent to learn to achieve the corresponding goal but also serve as a metric for evaluating whether the agent has achieved the goal. This metric is continuously monitored during the interactions of the CRL agent with the environment. When this average exceeds a predefined threshold, it is taken as an indication that the agent has successfully achieved the goals. The iterative process of high-level policy formulation, validation, and improvement draws inspiration from the concept of Reflexion Shinn et al. (2023). In instances where the agent fails to meet the threshold, indicating that the goal has not been met, the LLM is prompted to adjust the high-level policy. This adjustment is informed by feedback from the agent’s performance. The LLM first analyzes the feedback to pinpoint the reasons behind the shortfall in the previous policy. Then it leverages the powerful reasoning ability to hypothesize potential solutions. With this insight, the LLM formulates a new high-level policy aimed at overcoming the identified challenges.
3.2 Low-level Policy Learning
@@
The low-level policy learning is guided by the formulated high-level policy. The RL model learns the low-level policy using raw state input and total intrinsic reward based on the goal sequence . The agent’s intrinsic reward function is defined as:
@@
This reward structure provides a more granular view into the effectiveness of goal-related actions, encouraging the agent to learn a policy that is aligned with the goal more efficiently. Furthermore, the state-dependent design of this function improves the generalizability of high-level policies. Then, the internal rewards for each goal and the external reward from the environment will be textual and will be concatenated as the feedback. The total intrinsic reward is . Therefore, the aim of RL agent with goal sequence can be expressed as
@@ 3.3 Policy library
Recent studies Wang et al. (2023a); Zhang et al. (2023) have highlighted the potential of LLMs’ reasoning and planning ability in facilitating CL for agents, which involves continuous acquisition and update of skills Wang et al. (2023c). One method to address the core challenges in CL is to design a knowledge base Li et al. (2023) or a skill library Wang et al. (2023a). By using validated successful experiences, agents can not only reduce catastrophic forgetting but also rapidly enhance their capabilities.
Based on the above findings, Hi-Core maintains a policy library to store the policies that solve tasks successfully. The description of each high-level policy and the corresponding feedback is indexed by the embedding of the environment description of its corresponding task, which can be retrieved in similar tasks in the future. The policy library can be continuously extended throughout the life of the CRL agent. Given a new environment description , Hi-Core use embedding to retrieve the first similar environment descriptions and concatenate them with goal descriptions and corresponding feedback as experience . In addition, for each environment description , Hi-Core also stores the parameter of the low-level policy that has been learned. This can accelerate learning on similar tasks and further reduce catastrophic forgetting if storage space is sufficient. By storing both high-level and low-level learning policies in the policy library, the agent’s capabilities for diverse tasks can be enhanced over time.

### v3

# Multi-granularity Knowledge Transfer for Continual Reinforcement Learning

@@ Abstract
Continual reinforcement learning (CRL) empowers RL agents with the ability to learn a sequence of tasks, accumulating knowledge learned in the past and using the knowledge for problem-solving or future task learning. However, existing methods often focus on transferring fine-grained knowledge across similar tasks, which neglects the multi-granularity structure of human cognitive control, resulting in insufficient knowledge transfer across diverse tasks. To enhance coarse-grained knowledge transfer, we propose a novel framework called MT-Core (as shorthand for Multi-granularity knowledge Transfer for Continual reinforcement learning). MT-Core has a key characteristic of multi-granularity policy learning: 1) a coarse-grained policy formulation for utilizing the powerful reasoning ability of the large language model (LLM) to set goals, and 2) a fine-grained policy learning through RL which is oriented by the goals. We also construct a new policy library (knowledge base) to store policies that can be retrieved for multi-granularity knowledge transfer. Experimental results demonstrate the superiority of the proposed MT-Core in handling diverse CRL tasks versus popular baselines.
@@ 1 Introduction
Reinforcement learning (RL) is a powerful paradigm in artificial intelligence (AI) that enables agents to learn optimal behaviors through interactions with environments. The capacity to continuously adapt and learn from new tasks without starting from scratch is a defining characteristic of human learning Kudithipudi et al. (2022) and a desirable trait for RL agents deployed in the dynamic and unpredictable real world. To achieve this, the research of continual reinforcement learning (CRL, a.k.a. lifelong reinforcement learning) has emerged Rolnick et al. (2019); Kessler et al. (2022). It extends traditional RL by empowering agents with the ability to learn from a sequence of tasks, preserving knowledge from previous tasks, and using this knowledge to enhance learning efficiency and performance on future tasks. CRL aims to emulate the human capacity for lifelong learning and represents an ambitious field of research that tackles the challenges of long-term, real-world applications by addressing problems of diversity and non-stationarity Khetarpal et al. (2022).
@@
Despite progress in CRL, the field continues to face a critical challenge: insufficient transfer of knowledge across diverse tasks. Sufficient knowledge transfer is essential for a robust CRL system, yet many existing methods struggle to achieve this, leading to suboptimal performance and a failure to fully exploit the benefits of continual learning (CL) Wang et al. (2024c); Li et al. (2024). A primary reason for this shortfall is the neglect of the sophisticated multi-granularity (hierarchical) structure of human cognitive control, which is primarily located in the prefrontal cortex and is complemented by the cerebellum for precise action control, enabling coarse-grained cognition for envisioning the future and planning D’Mello et al. (2020); Friedman and Robbins (2022). This structure enhances humans’ ability to tackle complex tasks, and crucially, transfer skills and coarse-grained knowledge across significantly diverse tasks. In contrast, existing CRL methods often focus on transferring fine-grained knowledge, such as sharing the policy network (left side of Figure 1). This limits them to sequences of highly similar tasks with slight variations only in goals or parameters Kessler et al. (2022); Gaya et al. (2023), and results in ineffective transfer.
@@
To take advantage of abstract similarities between diverse tasks, it is essential to represent coarse-grained knowledge effectively and then employ robust models capable of extracting and transferring this knowledge. Human language, with its inherent abstraction, is a natural fit for representing coarse-grained knowledge. Furthermore, recent breakthroughs in large-scale language research have demonstrated the impressive reasoning and in-context learning abilities of the large language model (LLM) Zhao et al. (2023); Xi et al. (2025); Wang et al. (2024b). Based on these insights, we propose a new framework named MT-Core (Multi-granularity knowledge Transfer for Continual reinforcement learning), which integrates the powerful capabilities of an LLM into the CRL paradigm to enhance coarse-grained knowledge transfer and eventually improve learning performance.
@@
The right side of Figure 1 illustrates the idea of MT-Core. MT-Core is structured in two layers: the coarse-grained policy formulation and the fine-grained policy learning. At the coarse-grained, MT-Core utilizes the powerful reasoning ability of a LLM to formulate a coarse-grained policy, which is represented as a temporally extended sequence of goals. Each goal consists of a descriptive component and an intrinsic reward function that outlines the intermediate states the agent needs to reach. The coarse-grained policy is evaluated based on feedback from the fine-grained learning process, allowing the LLM to refine its policy. Following this process, the coarse-grained policy along with the corresponding information is stored in a policy library. Upon encountering a new task, MT-Core can retrieve this policy from the library as a context of the LLM for enhancing coarse-grained knowledge transfer. At the fine-grained, MT-Core employs goal-oriented RL to learn a policy that is guided by the formulated coarse-grained policy. The verified policy can be stored in the policy library to reduce the risk of catastrophic forgetting when encountering old tasks. Specifically, the coarse-grained policy is represented as text, while the fine-grained policy is represented as the policy network. The experimental results demonstrate that MT-Core significantly outperforms popular baselines. The main contributions of the work are as follows:
We investigate the knowledge transfer problem in CRL and find that the multi-granularity structure of human cognitive control is crucial for effective knowledge transfer across diverse tasks.
@@
We proposed MT-Core, a novel CRL framework that leverages multi-granularity knowledge transfer, which is the first work to integrate the powerful reasoning ability of LLM into the CRL paradigm, facilitating knowledge transfer across diverse tasks.
@@
Extensive experiments in MiniGrid provide empirical evidence of MT-Core’s effectiveness.
@@ 2.1 Preliminaries
The continual reinforcement learning process can be formulated as a series of related Markov decision processes (MDPs) . Each MDP represents a different task or problem instance that an agent needs to solve over its lifetime. Here, and denote the state and action space of task , respectively, while is the transition probability function, and is the reward function. At each time step, the learning agent perceives the current state and selects an action according to its policy with parameters . The agent then transitions to the next state and receives a reward . The target of an agent on the task is to maximize the expected return , where is the discount factor, and is the horizon.
@@ 2.2.1 Continual Reinforcement Learning
Another issue in CRL is transfer learning, which is crucial for efficient policy adaptation. Naive approaches, like fine-tuning, that train a single model on each new task provide good scalability and transferability but suffer from catastrophic forgetting. Regularization-based methods, such as EWC Kirkpatrick et al. (2017); Wang et al. (2024c), have been proposed to prevent this side effect, but often reduce plasticity. Some architectural innovations have been proposed to balance the trade-off between plasticity and stability Rusu et al. (2016); Berseth et al. (2022). Furthermore, methods like OWL Kessler et al. (2022) and MAXQINIT Abel et al. (2018) leverage policy factorization and value function transfer, respectively, for improved learning.
Most existing methods perform well when applied to sequences of tasks that exhibit high environmental similarity, such as tasks where only specific parameters within the environment are altered or the objectives within the same environment are different. However, their effectiveness is greatly diminished when dealing with a sequence of diverse tasks. Our proposed framework aims to overcome this limitation by leveraging multi-granularity knowledge transfer.
@@ 2.2.2 Reinforcement Learning With the LLM
Recent advancements have combined the LLM with reinforcement learning to address reinforcement learning challenges such as sample efficiency and generalization. These integrations utilize LLM’s vast knowledge and reasoning ability to improve agents’ performance. For instance, studies have developed methods for efficient agent-LLM interactions Hu et al. (2024) and unified agent foundations with language models for better experience reuse Palo et al. (2023a).
Frameworks that merge LLMs with RL have also emerged, such as LAMP, which pretrains RL agents with Vision-Language Models Adeniji et al. (2023), and HiP, which uses foundation models for long-horizon tasks Ajay et al. (2023). In addition, the RAFA framework offers a principled approach with guarantees to optimize reasoning and actions Liu et al. (2023). Research into LLM-guided skill learning and reward shaping is also gaining traction. The BOSS method Zhang et al. (2023) and the work on LLM-based reward structures for robotics Yu et al. (2023) exemplify this trend, with studies like Kwon et al. (2023) focusing on LLM-driven reward design. Furthermore, VLA combines coarse-grained visual-language policies and fine-grained reinforcement learning policies, demonstrating significant zero-shot transfer performance in the real world Yang et al. (2024).
CRL diverges from the traditional RL paradigm by focusing on complex and dynamic environments where the agent constantly encounters a sequence of tasks. Unlike other research on LLM support for RL, which mainly deals with static tasks or single tasks, our work integrates LLM’s powerful reasoning ability into the CRL paradigm for the first time. This distinction is crucial, as CRL poses unique challenges such as the necessity for high-level knowledge transfer across diverse tasks. Existing RL methods with LLM cannot address these challenges directly. Our framework’s novelty lies in its utilization of the LLM and the policy library not only to generate high-level policies that guide the learning process but also to facilitate high-level knowledge transfer between tasks, thus improving the agent’s adaptability.
@@ 3 Method
In this section, we first clarify the research problem and the overall framework. Then we elaborate on two layers of MT-Core: (1) the coarse-grained policy formulation, and (2) the fine-grained policy learning. Finally, we describe the policy library in MT-Core.
Problem Formalization.
@@
Our goal is to use the LLM’s powerful reasoning ability to enhance the capability of CRL agents. We formalize the CRL problem in the goal-oriented RL framework Chen and Luo (2022); Colas et al. (2022). The agent perceives not only the current state of the environment but is also provided with a specific goal , where represents the space of possible goals. For task and its corresponding goal sequence , the agent’s objective is to learn a policy that effectively maps each state to an appropriate action . The aim of the agent is to maximize the expected cumulative reward over time, guided by the sum of the extrinsic reward and the intrinsic rewards Dilokthanakul et al. (2019), where represents the next state in the task . The former is determined by the environment and is usually sparse. The latter is typically defined so that the agent receives a reward if it achieves the given goal.
Overview.
@@
MT-Core is illustrated in Figure 2. For simplicity, we have omitted the task-indexed superscript in the description. The framework orchestrates the interaction between the CRL agent and the environment simulator for each task in the task sequence, aligning with standard RL protocols. For each task, the corresponding environment simulator provides the current state to the agent at each timestep, which then selects an action to perform, resulting in a reward and the next state. Internally, the CRL agent is structured in two layers: the coarse-grained policy formulation and the fine-grained policy learning. Additionally, a policy library is maintained for the storage and retrieval of policies.
3.1 Coarse-Grained Policy Formulation
@@
To integrate LLM into the goal-oriented RL framework, we conceptualize LLM as a function , where is the set of all prompts, is the set of textual descriptions of the environment, and is the set of textual descriptions of the sequence of goals. The description of the environment can be automatically obtained through a hard-coded reporter Dasgupta et al. (2022) or a vision-language model Palo et al. (2023b); Du et al. (2023). The prompt is a concatenation of two parts: and , where is the system prompt and is the experience retrieved from the policy library. Description of each coarse-grained policy, that is, textual output is mapped to the corresponding goal sequence through a defined parsing function. In order to improve the stability of LLM’s output, we structure the goal into a tuple: , where is the textual description of the goal state, is a verification function , is a intrinsic reward value of the goal. Optional validation functions are provided in .
@@
The intrinsic reward can not only guide the agent to learn to achieve the corresponding goal but also serve as a metric for evaluating whether the agent has achieved the goal. This metric is continuously monitored during the interactions of the CRL agent with the environment. When this average exceeds a predefined threshold, it is taken as an indication that the agent has successfully achieved the goals. Consequently, the LLM shifts its focus to the next goal in the sequence, and the process repeats. The iterative process of coarse-grained policy formulation, validation, and improvement draws inspiration from the concept of Reflexion Shinn et al. (2023). In instances where the agent fails to meet the threshold, indicating that the goal has not been met, the LLM is prompted to adjust the coarse-grained policy. This adjustment is informed by feedback from the agent’s performance. The LLM is first required to analyze the feedback to pinpoint the reasons behind the shortfall in the previous policy. With powerful reasoning ability, the LLM then formulates a new coarse-grained policy aimed at overcoming the identified challenges.
3.2 Fine-Grained Policy Learning
@@
Fine-grained policy learning is guided by the formulated coarse-grained policy. The RL model learns the fine-grained policy using raw state input and total intrinsic reward based on the goal sequence . The agent’s intrinsic reward function is defined as:
@@
This reward structure provides a more granular view into the effectiveness of goal-related actions, encouraging the agent to learn a policy that is aligned with the goal more efficiently. Furthermore, the state-dependent design of this function improves the generalizability of coarse-grained policies. Then, the internal rewards for each goal and the external reward from the environment will be textual and will be concatenated as the feedback. The total intrinsic reward is . Therefore, the aim of RL agent with goal sequence can be expressed as
where is the parameter of fine-grained policy.
@@ 3.3 Policy Library
Recent studies Wang et al. (2024a); Zhang et al. (2023) have highlighted the potential of LLMs’ reasoning and planning ability in facilitating continual learning (CL) for agents, which involves continuous acquisition and update of skills Wang et al. (2024c). Two core challenges in continual learning are catastrophic forgetting and knowledge transfer Hadsell et al. (2020); McCloskey and Cohen (1989). One method to address the these challenges is to design a knowledge base Li et al. (2023) or a skill library to store knowledge or skillsWang et al. (2024a). By using validated successful experiences, agents can not only reduce catastrophic forgetting but also rapidly enhance their capabilities.
Based on the above findings, MT-Core maintains a policy library to store the policies that solve tasks successfully. The description of each coarse-grained policy and the corresponding feedback is indexed by the embedding of the environment description of its corresponding task, which can be retrieved in similar tasks in the future. The policy library can be continuously extended throughout the life of the CRL agent. Given a new environment description , MT-Core can use embedding to retrieve the first similar environment descriptions and concatenate them with goal descriptions and corresponding feedback as experience . Specific policies for new tasks can be learned under the guidance of coarse-grained goals. Specifically, We adopt a vector similarity search approach, with cosine similarity as the metric. When the number of tasks is small, simple storage strategies can also be used for the policy library. In addition, for each environment description , MT-Core also stores the parameter of the fine-grained policy that has been learned. This can accelerate learning on similar tasks and further reduce catastrophic forgetting if storage space is sufficient. Even if the storage space is strict, only storing coarse-grained policies (text) can perform knowledge transfer with small memory usage. By storing both coarse-grained and fine-grained learning policies in the policy library, the agent’s capabilities for diverse tasks can be enhanced over time.